• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Olson

    Log entry time 08:02:40 on May19,2007

    Entry number 202660

    keyword=Owl Shift Summary

    Owl Shift Summary:

    M.Olson, B.Hahn

    0021: MCC trying to locate and reboot ioc for Hall A BPM's (iocha10)

    0055: BPM readouts back, asked MCC for a harp scan at 5uA (BeO)

    0115: MCC needs to cycle a quad and rerun the harp scan

    0130: Harp scan (190 um): Starting to work on raster and beam center.
    0158: Raster established at 2.3mm x 2.3mm, working on beam spot
    0204: Beam spot within spec, asked MCC to mask target motion.

    Runs 3610-3613 (JUNK)
    Runs 3614-3618 (Raster and spot calibration)
    Run 3619 (JUNK)
    Run 3620 (He3 production, raster at ~2mm x ~2mm)

    0245: Requested raster increased by 15% (now 2.3mm x 2.3mm on spot++)

    0247: Run 3621 (He3 production)
    0317: End run 3621 on unstable beam. Request reduction to 90uA

    Good stable beam over last three hours...

    0605: Requested 95uA from MCC...to see if it will remain stable.
    0615: It won't remain stable at 95uA...request return to 90uA

    0642: MCC called, going to try for 105uA
    0700: Running at 105uA...reasonably stable
    0745: Good stability at 105uA

    ***Production Runs***
    3620 (2x2mm raster, 105uA, 366k)
    3621 (2.3x2.3mm raster, 105uA, 0.98M, ended on unstable beam)
    3622-3625 (1.25M events, 90uA)
    3626-3627 (1.25M events, 105uA)

    0801: Run 3628 in progress at end of shift.